MICROWAVE SPECTRA AND STRUCTURE OF CF_{3}I⋯CO

\underline{S. L. STEPHENS}, N. R. WALKER AND A. C. LEGON, School of Chemistry, University of Bristol, Bristol, BS8 1TS, U.K.

A Chirped Pulse Fourier transform Microwave spectrometer has been used to measure the pure rotational spectrum of CF_{3}I⋯CO. This complex is generated by supersonic expansion of a gas sample containing a small percentage of CF_{3}I, and CO in argon. The rotational constant B_{0}, centrifugal distortion constants, Δ_{J} and Δ_{JK} , and nuclear quadrupole coupling constant for iodine, χ_{aa}(I), have been determined for each of CF_{3}I⋯^{12}C^{16}O, CF_{3}I⋯^{13}C^{16}O and CF_{3}I⋯ ^{12}C^{18}O allowing determination of the distance between the two sub-units. The complex is a prolate symmetric top with C_{3v} symmetry.